Find the dimension of Planck constant h from the equation E=hv ,Where E is the energy and v is the frequency

Grade:12

3 Answers

Apoorva Arora IIT Roorkee
9 years ago
The Planck’s constant (h) can be defined as a proportionality constant that relates the energy (E) of a photon to the frequency (ν) of its associated electromagnetic wave.
Mathematically,
Planck’s Constant (h) = Energy(E) / frequency (ν).
Dimensional Formula of energy = M1L2T-2.
Dimensional Formula of frequency = M0L0T-1.
Putting these values in above equation we get,
Dimensional Formula of Planck’s constant = M1L2T-1.
SI unit of Planck’s constant is joule-seconds (j-s).
Anjali kumari
5 years ago
It should be ML^2T^-1. As planks constant,h = E/neu I.e. frequency. Dil will aap aap O do ciao ph artho spark zo pap Ozark.
As energy,E=ML^2T^-2 and the dimension of frequency, v=T^-1.ThereforeE=hv. So, 
ML^2T^-2=h T^-1.Hence, h=ML^2T^-1.
Ajeet Tiwari
3 years ago
hello student


The Planck’s constant (h) can be defined as a proportionality constant that relates the energy (E) of a photon to the frequency (ν) of its associated electromagnetic wave. Mathematically, Planck’s Constant (h) = Energy(E) / frequency (ν). Dimensional Formula of energy = M1L2T-2. Dimensional Formula of frequency = M0L0T-1. Putting these values in above equation we get, Dimensional Formula of Planck’s constant = M1L2T-1. SI unit of Planck’s constant is joule-seconds (j-s).

hope it helps
thankyou